The first order reaction, SO2Cl2 -> SO2 + Cl2, has a rate constant of 0.17 h^-1. If the initial concentration of SO2Cl2 is 1.25 x 10^-3 M, how many seconds does it take for the concentration to drop to 0.31 x 10^-3 M?
Answer #1

Initial concnetration, a = 1.25 * 10-3 M

Final concentration, (a-x) = 0.31 * 10-3 M

Rate constant, k = 0.17 h-1

TIme, t = ?

The first order rate constant equation,

k = (2.303 / t) Loga/(a-x)

t = (2.303 / 0.17) * Log (1.25*10-3/0.31*10-3)

t = 8.20 h

t = 8.20 * 60 * 60 s = 29520 s